P(L)=(pi2EI)/L2

A vertical column buckles when too large a weight (P) is placed on it. the load that will buckle a column of length L is given by the formula above. E and I are constants that depend on the shape and material of the column. Sketch a neat graph of the function P(L).

Hint: Increment the P-axis with increments of pi2EI since we don't know what E and I are. Before graphing try small values of L and then slowly increase or decrease L.
Answer #1

First, i drew the graph of y = 1/x^2, and then related it to P = C/L^2. I then showed how The graph would vary by changing the values of C. As C is a material property, we have a wide range of C. I took 5 values of 0.001*C,( C = 1,50,100,300,700,1000) and plotted it in python for accurate results